HUBBLE SPACE TELESCOPE DAILY REPORT #936 PERIOD COVERED: 0800 EST Daily Status Report as of 207/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of PC 4809 (Cycle 2: WF/PC PSF Calibration Proposal) The Planetary Camera was used to make a point spread function calibration image of standard star HD151406. The OSS monitored the observation and assessed the data quality. No problems were reported. 1.2 Completed FGS 2565 (Linking the HIPPARCOS to the Extragalactic Reference Frame Part 1 of 6) The Fine Guidance Sensors were used to make two sets of astrometric observations of objects 0109+224 and object 142REF. These objects lie near the apparent position of extragalactic objects. These objects will also be observed with the HIPPARCOS satellite. The intent is to tie the HIPPARCOS system reference frame to the extragalactic reference frame. The observations were executed as scheduled, and no problems were reported. 1.3 Completed Two Sets of FOS/RD 4953 (QSO Absorption System Snapshot Survey - Part 3 of 3: Cycle 3 High) The Faint Object Spectrograph was used to make observations of quasar 1612+2611 to look for absorption lines at low redshift, and also emission lines in the QSO's continuum. The observations were executed as scheduled, and all the observations were reported to be okay. 1.4 Completed Four Sets of WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Deltas Flats) The Wide Field Camera was used to make a total of four internal flat calibrations images. All four images were reported to be normal. 1.5 Completed GHRS/2 5012 (Physical and Chemical Processes in Dense Interstellar Clouds) The High Resolution Spectrograph was used to make observations of star HD154368. This target star is behind a translucent interstellar cloud which is dense enough for the formation of molecules, but still passes some star light. The proposers are interested in the absorption lines in the stellar spectrum. The observations were complete as planned, and no problems were reported. 1.6 Completed Two Sets of GHRS/2 3941 (Physical Parameters in the Local Interstellar Medium - Cycle 2) The High Resolution Spectrograph was used to make observations of the star HD209952 (aka, Alpha Gruis) to observe the weak lines from neutral atomic species in the local interstellar medium. The baseline guide star acquisition made fine lock (2 star!). The on-board spiral target acquisition located the target on the third map. The resulting spectral observations showed plenty of signal. There is apparently a new dead diode in channel 471. 1.7 Completed Two Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Narrow. 2) The Wide Field Camera was used to make Earth flat calibration images to be used for flat fielding of WFC observations. The observations used the narrow band filters, both with and without neutral density filters. The observations were all received. No problems were reported. 1.8 Completed Eight Sets of PC 4714 (WF/PC Cycle 2 Calibration: PC Internal Delta Flats) The Faint Object Camera was used to make two internal flat calibration images which will be used for the restoration of Earth flat images, which were altered due to changes in quantum efficiency which occurred after decontamination operations. Both observations were received and displayed in the OSS. They were reported to look okay. 1.9 Completed FOC/96 3870 (Imaging of M-31 Group Galaxies - Cycle 2) The Faint Object Camera (f/96) was used to make seven images of the object NGC-224 which is located near galaxy M-31 (The Andromeda Galaxy) to determine the population and gradients of this object. The observations were executed in fine lock. The images were received and displayed, and no problems were reported. 1.10 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.11 Completed FGS 5068 (Linking the HIPPARCOS to the Extragalactic Reference Frame Part 6 of 6) The Fine Guidance Sensor #3 was used to make two sets of astrometric observations of object NEWEGOD-62. This object lies near the apparent position of extragalactic objects. This object will also be observed with the HIPPARCOS satellite. The intent is to tie the HIPPARCOS system reference frame to the extragalactic reference frame. The observations were executed as scheduled, But there was a problem with the astrometry. Whenever FGS #3 looked at the target NEWEGOD-62 the stop flag, and the scan step limit exceeded flag went out of limits. 1.12 Completed FOS /BL 4328 (The Companions of Dwarf Carbon Stars: Cycle 3 Medium) The Faint Object Spectrograph was used to make spectral observations of the dwarf carbon star LHS1075 which is suspected of having a cool white dwarf companion. There were no mechanical or software problems with the peakups. However, the counts were very low, and nearly uniform in each group. The target is either too faint for the exposure time used, or it is nowhere near the field of view. 1.13 Completed Two Sets of FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) The Faint Object Camera was used to complete two sets of snapshot ultraviolet observations fo the nearby galaxy ESO-2312-4 and galaxy ESO-1115-3. The exposures lasted for ten minutes and used filter F220W. Both sets of observations were reported to be fine. 1.14 Completed GHRS/2 3935 (Sulfur Near IO - Cycle 2) The High Resolution Spectrograph was used to make a series of spectral observations of Jupiter's moon Io while it was both inside and outside of the plasma torus. This is to determine the mechanism which feeds sulfur and oxygen into the torus. The onboard target acquisition was successful. All observations were reported to look okay, although they were rather weak. 1.15 Completed FOC/96 4800 (The Stellar Content of Wolf- Rayet Galaxies (FOC Images)) The Faint Object Camera was used to make one spectral observations of the Wolf-Rayet galaxy TOL-89. This galaxy is a star bursting galaxy which exhibits broad He II emissions at 4686 A0 from stellar sources. The observation was received and displayed. No anomalies were noted. 1.16 Unscheduled Safemode Entry The HST entered software sunpoint safemode at approximately 03:24 A.M. EDT Monday 26, 1993. (See details in Operations Note 2.4). 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 19 Successful : 19 Scheduled Reacquisitions : 18 Successful : 18 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 38 Successful : 38 2.3 Significant HSTARs: None 2.4 Operations Note: SOFTWARE SUNPOINT SAFEMODE ------------------------- The HST entered software sunpoint safemode at approximately 03:24 A.M. EDT Monday 26, 1993. The Spacecraft on-board computer detected a difference between the expected and actual values for the solar array wings during a scheduled wing maneuver of 30 degrees. The present limit for difference is 8.0 degrees and the safemode was activated when this limit was exceeded by the -V2 wing. A previous solar array slew of 30 degrees approximately 24 hours earlier had executed normally. Initial analysis indicates a nominal safemode entry with all subsystems in a safe state. The payload went to the expected "soft safing" condition. The vehicle is power positive and under nominal safemode ground control by STOCC controllers. Data relevant to the safemode entry is on the on-board tape recorder since the event occurred during a ZOE. The earliest time this data can be dumped to a ground station is about 5:15 P.M. toady. The data will be dumped to Madrid and then forwarded to GSFC. Data analysis will begin as soon as the data is available at GSFC. Representatives from ESA are at GSFC and are supporting the investigation. Steps are under way to get the spacecraft operational on a Health and Safety Load to provide more positive and efficient ground operations control, and to facilitate the return to science operations when the problem is better understood and resolved. The H&S load will be developed today and is expected to start executing later tonight. Plans for returning to science operations, currently on hold, will be developed after the problem is better understood. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC 4599 (Optical Morphology of Radio Quiet, Cluster Galaxies at Redshift Z = 3.4) 3.2 Complete FOS/BL 5050 (Cycle 3 Internal Wavelenght Calibration) 3.3 Complete Two Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.4 Uplink Updated Ephemeris Tables () 3.5 Complete FOS/RD 4772 (Black Holes, Stellar Dynamics and Populations in the Nuclei) 3.6 Complete WFC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-